Status Report

NASA Space Station On-Orbit Status 23 Jun 2003

By SpaceRef Editor
June 23, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 23 Jun 2003

All ISS systems continue to function nominally, except as noted previously or below.  Week 8 of Increment 7 is underway.

During a Russian ground site pass on Daily Orbit 3 (DO3), CDR Yuri Malenchenko performed his second session with the Russian MedOps cardio experiment MO-1 (Study of the Bioelectric Activity of the Heart at Rest), with FE/SO Ed Lu assisting as Crew Medical Officer (CMO).  [First time done: 5/27. During the test, transmittal of the Gamma-1M ECG (electrocardiogram) signal to the ground was required for about 5-6 minutes.]

In a late change to his timeline for today, Malenchenko spent several hours troubleshooting the cable network for the +X Klest140st-M television camera of the Progress 10P docked at the Service Module (SM) aft end.  The activity was supported by TsUP/Moscow during DO4 with a test of the camera via the Progress Kvant radio command system.  [Using the MMTs-01 multimeter, Yuri took Ohm (resistance) readings to check electrical connectivities and looked for possible mechanical damage to insulation of cable branch points to demated connectors.  The Klest system is part of the Progress TORU approach and docking system.]

The crew started this week’s maintenance work on the RED (resistive exercise device) in a three-hour IFM (in-flight maintenance) operation, swapping the Flexpack in RED canister #1009 with a new unit brought up on 11P.  The second canister, #1010, will receive its new Flexpack on 6/25 (Wednesday).  [After the new Flexpacks have been installed, the old units will be inspected to ascertain how many good packs are available for use as spares when the current Flexpacks fail.  The inspection is scheduled as a task list item starting on 6/25 and hard scheduled during Week 9.  The replaced units from 1009 and 1010 will be used in canisters 1001 and 1003 in a separate IFM once these cans are considered failed.]

Yuri Malenchenko continued the latest round of periodic preventive maintenance of Russian segment ventilation systems, today cleaning the Group A fans and grilles in the SM (last time done: 5/21),

Ed Lu performed an inspection of the VOA (volatile organics analyzer) which has exhibited some offnominal reboots during sampling and data dump, with flow rate being lower than normal ever since the last remote sampling.  [The information from the inspection will help engineers to troubleshoot the low flow rate and provide possible causes of the reboots.]

Later, the ground commanded a VOA (volatile organics analyzer) calibration run, lasting about 3.5 hrs., then beginning another 3.5 hrs. of air sample collection. 

The Science Officer also collected air samples with the GSC (grab sample container), then deployed two passive FMK (formaldehyde monitoring kit) sampling badges in the Lab and SM, to catch any atmospheric formaldehyde on a sampling substrate for analysis in JSC lab facilities (last time done: 5/26/03).

Malenchenko set up the equipment for tomorrow’s repeat of the downlink of the video footage of the geophysical GFI-1 Relaksatsiya (“relaxation”) experiment with Soyuz 5S deorbit burn plumes, Earth limb and Turkish landmass views tomorrow, except for the DVCAM battery, for which overnight charging will be initiated later today.  [The previous downlink attempt of the footage on 5/20 was not completed satisfactorily since it did not contain the desired signal.]

Ed Lu performed the periodic functional closure test of the Vozdukh CO2 removal system’s emergency vacuum valves (AVK, last time done: 5/21).  [The AVKs are critical because they close the Vozdukh’s vacuum access lines in the event of a malfunction in the regular vacuum valves (BVK) or a depressurization in the Vozdukh valve panel (BOA).  Access to vacuum is required to vent carbon dioxide during the regeneration of the absorbent cartridges (PP).  During nominal operation, the AVK valves remain open.]

Lu also reviewed the video test pattern downlink procedure of the HRF (human research facility) Ultrasound apparatus in a computer-based training session (CBT).  This was in preparation for tomorrow’s scheduled HRF/Space Medicine joint Ultrasound  activity.

The CDR performed the regular periodic inspection of the BRPK-1 air/liquid condensate separator in the SM.  Yuri also completed his regular daily monitoring/servicing of the BIO-5 Rasteniya-2 (“Plants-2”) zero-G plant growth experiment, checking water level in the Lada-2 greenhouse water container and seed status.

Ed conducted the regular status checkup of the autonomously running Increment 7 Lab payloads (PCG-STES01, SAMS, MAMS).

The first data take and downlink of the NGL (Next Generation Laptop) machine last Friday (6/20) was successful.  This means that the NGFL is now “off and running” on its three-week zero-G shakedown test.  [The NGL, of Notebook class, has a 2GHz Pentium-4 processor, a 60 GB hard disk drive (HDD), 1 GB RAM memory, a 2-hr. Li-Ion battery and a mass of less than 8 lbs.]

U.S. control moment gyroscopes (CMGs) #1 and #2 each suffered a brief loss-of-comm event with the GNC MDM (guidance, navigation & control computer).  Both were recovered quickly with auto-retry and manual power cycle; they are now running nominally.

After last week’s (6/17) thruster tests on Progress 11P from the SM computers via the newly installed US-21 matching unit, a Russian commission is now looking into possible contamination effects on the SM solar cell arrays due to thruster plumes from the cargo ship, which is docked radially at the DC-1 docking compartment in relative proximity to the array wings.
Today’s CEO (crew earth observation) targets, limited in the current XPOP attitude by flight rule constraints on the use of the Lab nadir/science window, and including the targets of the Lewis & Clark 200-year memorial locations, were Three Gorges Dam, Yangtze River (daytime cumulus were likely at the time of the ISS pass, but good views of the Yangtze in the vicinity of the dam lay either side of track), Mekong River Delta (the lower Mekong River is undergoing significant changes as dams and reservoirs are completed upstream. Where clouds permit, the crew was to try for views either side of track noting the land use and water color), Gulf of Maine plankton (poor weather limited visibility to the south, but crew was advised to look left of track for open water and possible plankton blooms between the coast of Maine and southwestern Nova Scotia), Mouth of Bad River, Fort Pierre (LEWIS & CLARK SITE:  Two passes today, of which the second was perhaps the better.  The Bad River is a small, meandering stream that enters the Missouri River from the southwest, just opposite from Pierre, SD.  Here the expedition first encountered tension and potential hostility in meetings with the Lakota Indians.  Looking right of track for the city of Pierre, on the Missouri, just south of the dam for Lake Oahe), and Missouri River Sites (LEWIS & CLARK SITE(S):  On May 14, 1804, the expedition started up the Missouri river in a 55-ft keelboat and two smaller craft from near St Louis.  Looking left of track for the large modern city and especially on the north side where the Missouri and the Mississippi Rivers slowly meander away from one another).
CEO images can be viewed at the websites

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 12:40pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (18 amp mode).  Vozdukh CO2 scrubber is On (in Manual Mode 3/5).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 28.9; ppO2 (mmHg) — data invalid; ppCO2 (mmHg) — data invalid.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.9.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 22.0.
  • Node:  Pressure (mmHg) — 745.80; temperature (deg C) — 24.4 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 747.64; temperature (deg C) — 22.9; ppO2 (mmHg) — 172.4; ppCO2 (mmHg) — 3.9.
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 747.75; temperature (deg C) — 30.8; shell heater temp (deg C) — 25.5, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 25.5.
  • PMA-2:  Shell heater temp (deg C) — 13.8.

(n/a = data not available)

Propulsion System (PS):

  • Total propellant load available: 3803 kg (8384 lb) as of 6/19  [SM(774) + FGB(2447) + Progress M(182) +Progress M-1(400)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in AutoTrack.
  • SM batteries:  Battery #4 is disconnected (failed 6/16); all other batteries (7) are in “Partial Charge” mode.
  • FGB batteries:  Battery #4 is in “Cycle” mode; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-1 is operating; INT-2 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is Off; PL-1 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist). Management.
  • Solar Beta angle:  -26.9 deg (magnitude decreasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally [IAC-1 (internal audio controller #1) being analyzed after self-test error.  IAC-2 is prime.]
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 3:22am EDT [= epoch]):

  • Mean altitude — 388.2 km
  • Apogee — 392.9 km
  • Perigee — 383.5km
  • Period — 92.32 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0006931
  • Orbits per 24-hr. day — 15.60
  • Solar Beta Angle — -26.9 deg (magnitude decreasing)
  • Mean altitude loss in last 24 hours — 60 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 26196
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.